# Few-shot Species Range Estimation

### v1

@@ B.4 Ablating Location Encoder
In Figure A4 we vary the number of datapoints used to pretrain the SINR encoder used in FS-SINR. For both FS-SINR and the SINR baseline, we generally observe that more data is better, and for SINR approaches we see that pretraining the encoder is much better than randomly initializing it. We also show results for a SINR model trained on evaluation species as well as train species. As we saw in Table 1 for FS-SINR, the impact of training the location encoder with evaluation species is small.
@@
In Figure A5 we also investigate the impact of changing the location encoder entirely. We see that replacing our SINR location encoder with a pretrained and frozen “SATCLIP” location encoder (Klemmer et al., 2023) significantly harms performance. This may be due to this model being frozen and trained on tasks that do not completely match ours. In comparison, a randomly initialized and untrained SINR backbone performs almost identically well as one that has seen a small amount of training data (10 examples per-species in the train set). We also investigate replacing the learned location encoder $f()$ with a simple form of Fourier feature encoding (Tancik et al., 2020) for encoding inputs to $m_{\bm{\psi}}()$. In this setting, a pretrained and finetuned SINR type location encoder $f()$ is still used to encode locations $x$ to determine the probability of presence of species $j$ via the inner product between $\bm{w}_{j}$ and $f(x)$. However, $f()$ is not used to encode $\mathcal{C}^{t}$ before it is passed to the transformer $m_{\bm{\psi}}()$. Using these two different encoders performs increasingly poorly as the amount of context information increases.
@@
Figure A4: Impact of Location Encoder Training. Here we evaluate the performance of SINR and FS-SINR models when the size of the training dataset for the SINR backbone is varied. Results for FS-SINR models are shown with standard deviations from three runs (left), and without (right) for clarity. Evaluation on FS-SINR is performed with “Range Text”, while SINR can only make use of location data. “1000”, “100”, and “10” represent the maximum number of examples per class the SINR backbone was trained on. “SINR (rand_init)” is initialized with random weights and is not trained. “(trained on eval species)” means the model was trained on all training and evaluation species.
@@
Figure A5: Impact of location encoder. Here we evaluate the performance of FS-SINR style models with different location encoders. Results are shown with standard deviations from three runs (left), and without (right) for clarity. Evaluation is performed with “Range Text” on the IUCN dataset. “1000”, “100”, “10” represent the maximum number of examples per class the SINR backbone was trained on. “(Frozen)” indicates that the location encoder parameters were not updated during FS-SINR training. “FS-SATCLIP” replaces the SINR location encoder with a pretrained, frozen location encoder from Klemmer et al. (2023). “FS-SINR (Fourier Location Encoder)” uses the simple Fourier feature encoding (Tancik et al., 2020) used in Mildenhall et al. (2021) to match the 256 dimensional outputs of the SINR location encoders. These outputs are used directly as inputs to the transformer encoder. After a species token is produced in this way, it is attached to a pretrained and finetuned SINR backbone to produce a range.
B.5 Ablating Training Data
In Figure A6 we vary the number of examples per-species that are provided during training. The impact of this is fairly small, with models trained on an intermediate amount of data performing best. It is worth noting, that not all species in the training dataset have as many as 1000 observations. We find that the a model trained on only 10 examples per-species performs significantly worse.
@@ B.5 Ablating Training Data
Figure A6: Impact of training data. Here we evaluate FS-SINR models trained with different amounts of data. Results are shown with standard deviations from three runs (left), and without (right) for clarity. Evaluation is performed with “Range Text” on the IUCN dataset. The labels show the maximum number of examples per-species that FS-SINR is trained on. We see that training on an intermediate amount of training data leads to best performance.
B.6 Ablating FS-SINR Architecture
In Figure A7 we vary the underlying FS-SINR architecture. Removing different components has a small effect on model performance, with the removal of the species decoder actually improving results when range text is provided. However, as several ablations perform very similarly, it is difficult to tease out the how much of this effect is due to variance. It is clear however that removing the learnable token type embeddings causes the model to completely fail to learn during training. In Figure A8 we show further ablations based around removing the learned location encoder for inputs to the transformer and replacing it with the simple Fourier feature encoding also seen in Figure A5. When this is removed, other ablations seem to further harm performance, though results for these ablations vary a lot between runs.
@@ B.6 Ablating FS-SINR Architecture
Figure A7: Ablating model architecture components. Here we evaluate the performance of FS-SINR style models as we ablate various design choices. Results are shown with standard deviations from three runs (left), and without (right) for clarity. Evaluation is performed with “Range Text” on the IUCN dataset. We see small changes in performance when removing the register token and the species decoder. However removing the learned token type embeddings has a large impact.
@@
Figure A8: Further ablating model components. Here we evaluate the performance of FS-SINR style models as we ablate more components. Results are shown with standard deviations from three runs (left), and without (right) for clarity. Evaluation is performed with “Range Text” on the IUCN dataset. “FF” indicates that the model does not use a SINR backbone to encode location inputs to the transformer encoder. Instead, a simple Fourier feature encoding (Tancik et al., 2020) used in Mildenhall et al. (2021) is used to increase the dimensionality of location data to match the token dimension of the transformer encoder. These are used directly as inputs to the transformer encoder. After a species token is produced in this way, it is attached to a standard SINR backbone to produce a range. Removing the SINR backbone for encoding inputs to the transformer has a large impact on performance, especially when more context locations are supplied, and makes the model more sensitive to the impact of other ablations.
B.7 Taxonomic Understanding
Here we investigate the impact of providing FS-SINR with an understanding of the species taxonomy. For this we provide “Taxonomic Rank Text” (TRT) instead of the Wikipedia-based free-form descriptions of a species that are used for our standard FS-SINR approach. This text gives the taxonomy of the species in decreasing taxonomic rank, in the form “class order family genus species”, so for a dog we would give the text “Mammalia Carnivora Canidae Canis Familiaris”. During training, we select a rank uniformly at random and remove all ranks underneath that. We hope that this process will force the model to learn an understanding of the distributions of not only individual species, but also genera, families, etc.. This may be helpful when facing unseen species as knowledge of the genus or family may provide clues about where this species may be found. This is similar to the approach used by LD-SDM (Sastry et al., 2023).
In Table A2 we show zero-shot performance for FS-SINR models trained on TRT on the IUCN and S&T evaluation tasks. We see that as we provide additional taxonomic information zero-shot performance improves, though it is still much worse than using habitat or range text. This implies that the model has managed to develop some understanding of the distributions of genera etc. and can use this to help it map a novel species that shares higher order taxonomy with species in the training set. In Figure A9 we show few-shot results for FS-SINR models trained on TRT on the IUCN and SNT evaluation datasets. Zero-shot performance improvement with increasing taxonomic information is evident, but after very few provided locations this effect seems to disappear.
In Figure A10 we provide some qualitative zero-shot and few-shot results showing the impact of training on taxonomic text. We see that the model appears to narrow down on the correct range as more specific taxonomy is revealed to it, from predicting across the entire globe when just the class Aves is provided, to removing northern latitudes as the family Columbidae is added, and finally removing the new world when the genus is provided. This broadly matches the actual distribution of these taxonomic ranks. Note, the relationship between taxonomic hierarchy and species range is likely complex as many speciation events (i.e., when a species splits into two or more new ones) can be the result of physical geographic barriers separating populations over time.
Figure A9: Impact of training and evaluating with Taxonomic Rank Text. Here we evaluate FS-SINR models trained using different context information on the IUCN dataset (left), and the S&T dataset (right). “Class” indicates that only the taxonomic class of the species is provided as text during evaluation. “Order” indicates that the taxonomic class followed by the order is provided as a text string during evaluation, and so on, such that “Species” indicates that a text string in the format “class order family genus species” is provided during evaluation. Providing more specific taxonomic text increases zero-shot performance. This is also presented Table A2. However we see that even the full taxonomy does not provide as much signal as habitat and range text for zero-shot range mapping. These more detailed texts provide more useful information for zero-shot range mapping - either actually mentioning geographic locations in the case of range text, or allowing the model to narrow predictions down to areas with specific features such as mountains and forests in the case of habitat text. When a single context location is provided, the choice of taxonomy text no longer seems to impact performance at all. It is possible that training on these less informative tokens means the model learns to pay less “attention” to these text tokens compared to the Wikipedia-based text tokens usually used during training. This could explain why different rank taxonomy text tokens seemingly provide no benefit when any context locations are provided to the model.
Method Variant IUCN S&T
FS-SINR 0.05 0.18
FS-SINR HT 0.33 0.53
FS-SINR RT 0.52 0.64
FS-SINR Class 0.05 0.19
FS-SINR Order 0.06 0.20
FS-SINR Family 0.12 0.25
FS-SINR Genus 0.18 0.30
FS-SINR Species 0.21 0.34
Class
Aves
Family
Columbidae
Genus
Treron
Species
Phoenicopterus
[TRT Text]
[TRT Text] + 1 Context Location
Figure A10: Zero-shot and one-shot range estimation using Taxonomic Rank Text (TRT). Range predictions for the species Yellow-footed Green Pigeon from FS-SINR model trained on taxonomic rank text as in LD-SDM (Sastry et al., 2023), with expert-derived range inset. As seen in Figure A9 and Table A2, the text-based zero-shot predictions seem to more closely match the expert-derived range as more of the taxonomic rank text of the species is provided. Taxonomic rank text allows the model to somewhat localize predictions to areas where species sharing the provided taxonomy ranks are present in the training set. For example, Birds are globally distributed and we see the model attempt to output this in the zero-shot ‘Class’ visualization. Pigeons and Doves are not found in the extreme north and providing these ranks reduces predictions in these areas (and much of the northern hemisphere). The model mostly manages to identify that Green Pigeons are found only in Africa and parts of Asia. A single observation significantly contracts the predicted ranges, particularly when less taxonomic information is provided. Click on the species names to visit the iNaturalist page for that taxonomic rank to see the geographic distribution of observations of that taxa, which may resemble that in our training data.
Appendix C Additional Qualitative Results
In this section we provide additional qualitative results.
C.1 Qualitative Results
@@
As in LE-SINR (Hamilton et al., 2024), by jointly training on text and locations, FS-SINR is able to spatially ground abstract non-species concepts in a zero-shot manner. In Figure A11 we see some examples where different text concepts, that are very different from the species range or habitat text provided during training, are grounded in sensible locations on the map. In Figure A12 we compare models with and without text cues. As we increase the number of context locations, the two different models converge to more similar range predictions. In Figure A13 we provide another example similar to Figure 4 in the main paper. Here, we again fix the context location and show the impact of changing the text. We can see that different text prompts can result in quite different predicted ranges. In Figures A14 and A15 we visualize the model range predictions for two different species when richer habitat or range text is provided. We observe that the combination of text and context locations (here only location is provided) results in the best performance. Finally, in Figure A21 we show some additional zero-shot image only examples, where FS-SINR gets a single image from a held-out test species at inference time and attempts to generate a range map. Again, we observe some plausible range predictions even with such limited input data.
@@
In Figure A16 we visualize FS-SINR range predictions for the Yellow-footed Green Pigeon for models that have had different random initializations (i.e., different random seeds). We observe that there is a relatively large amount of variance in the outputs produced given the same input data. The same set of input context locations could represent many different possible output ranges, and thus being able to represent this variety is advantageous.
@@
[…]
(a) European Robin - Range Text
(b) European Robin - Habitat Text
(c) The European robin is found across Europe, east to Western Siberia and south to North Africa; it is sedentary in most of its range except the far north. It also occurs in the Atlantic islands as far west as the Central Group of the Azores and Madeira. It is a vagrant in Iceland and has been introduced to other regions, including North America and Australia, but these introductions were unsuccessful.
(d) The European robin inhabits a variety of habitats, including gardens, parks, woodlands, and forests. It prefers areas with dense vegetation and is often found near human settlements. It is also found in mountainous regions and can be seen in urban areas, such as cities and towns.
Figure A14: Using text descriptions. Here we illustrate the zero-shot (top row) and one-shot (bottom row) range estimations based on text descriptions for the European Robin, using ‘Range’ (left), and ‘Habitat’ (right) text, shown below the range estimates. Expert-derived range maps are shown inset in the top row.
(a) American Pika - Habitat Text
(b) American Pika - Range Text
(c) The American pika (Ochotona princeps) is found in the mountains of western North America, usually in boulder fields at or above the tree line, from central British Columbia and Alberta in Canada to the US states of Oregon, Washington, Idaho, Montana, Wyoming, Colorado, Utah, Nevada, California, and New Mexico.
(d) Pikas inhabit talus fields that are fringed by suitable vegetation in alpine areas. They also live in piles of broken rock. Sometimes, they live in man-made substrate such as mine tailings and piles of scrap lumber. Pikas usually have their den and nest sites below rock, around 20-100 cm (8-39 in) in diameter, but often sit on larger and more prominent rocks.
Figure A15: Using text descriptions. Here we illustrate the zero-shot (top row) and one-shot (bottom row) range estimations based on text descriptions for the American Pika, using ‘Range’ (left), and ‘Habitat’ (right) text. Expert-derived range maps are shown inset.
@@
C.2 Visualizing Embeddings
In Figure A17 we show Independent Component Analysis (ICA) projections of the location encoder features for FS-SINR, LE-SINR, and SINR. We encode all locations around the world into 256 dimensional representations by passing them through the location encoders from the different models, and we then reduce these to three dimensions and visualize them as RGB colors as in Cole et al. (2023). We also display this for an FS-SINR model trained on taxonomic rank text. Locations with similar colors should have similar location features and represent locations that the model thinks may share species. Across all models, higher frequency changes in location features are seen in areas where we have more training data. This can be seen particularly clearly by comparing the United States and Europe compared to central Asia or Africa.
(a) FS-SINR
(b) FS-SINR (TRT)
(c) LE-SINR
(d) SINR
Figure A17: Visualization of the learned features of different location encoders. Here we project high dimensional location features down to three dimensions using Independent Component Analysis.
C.3 Qualitative Comparisons
Here we present qualitative comparisons of the ranges produced by FS-SINR, LE-SINR, and SINR. In Figure A18 we show range estimates for the Brown-banded Watersnake, using range text for FS-SINR and LE-SINR approaches. In Figure A19 we show range estimates for the Brown-headed Honeyeater, using habitat text for FS-SINR and LE-SINR approaches. Finally in Figure A20 we show range estimates for the Crevice Swift, without providing text. Overall, SINR produces more diffuse ranges and requires more locations to narrow down the range. LE-SINR and FS-SINR appear to have very different zero-shot behaviors, with LE-SINR frequently seeming to predict presence in almost no locations at all, while FS-SINR tends to produce a zero-shot range that is too large.
@@ 20 Context
Figure A18: Comparing estimated ranges across models with context points and range text. Here we see zero-shot and few-shot range estimates produced by FS-SINR, LE-SINR, and SINR for the Brown-banded Watersnake, with expert-derived range inset. We provide range text to FS-SINR and LE-SINR as well as context locations, but SINR is not capable of accepting text and so we show a blank map for the zero-shot range estimate. We see that LE-SINR underestimates the range using only text, while FS-SINR overestimates it. SINR requires more location data than the other approaches to localize the range to South America.
@@
Figure A19: Comparing estimated ranges across models with context points and habitat text. Here we see zero-shot and few-shot range estimates produced by FS-SINR, LE-SINR, and SINR for the Brown-headed Honeyeater, with expert-derived range inset. We provide habitat text to FS-SINR and LE-SINR as well as context locations, but SINR is not capable of accepting text and so we show a blank map for the zero-shot range estimate. We again see LE-SINR underestimate the range using only text, while FS-SINR has very good zero-shot performance for this species. We see that SINR again requires more location data to narrow down the range and even after 20 locations the range is still significantly larger than the other models, extending into South Africa.

### v2

@@ D.4 Ablating Location Encoder
In Figure A14, we vary the number of datapoints used to pre-train the SINR encoder used in FS-SINR. For both FS-SINR and the SINR baseline, we generally observe that more data is better, and for SINR approaches we see that pretraining the encoder is much better than randomly initializing it. We also show results for a SINR model trained on evaluation species in addition to train species. As we saw in Table 1 for FS-SINR, the impact of training the location encoder with evaluation species is small.
@@
In Figure A15, we also investigate the impact of changing the location encoder entirely. We see that replacing our SINR location encoder with a pre-trained and frozen ‘SATCLIP’ location encoder (Klemmer et al., 2025) significantly harms performance. This may be due to this model being frozen and trained on tasks that do not completely match ours. In comparison, a randomly initialized and untrained SINR backbone performs almost identically well as one that has seen a small amount of training data (10 examples per-species in the train set). We also investigate replacing the learned location encoder $f()$ with a simple form of Fourier feature encoding (Tancik et al., 2020) to encode location inputs to the transformer $m_{\bm{\psi}}()$. In this setting, a pre-trained and fine-tuned SINR type location encoder $f()$ is still used to encode evaluation locations $x$ to determine the probability of presence of species $j$ via the inner product between the species embedding vector $\bm{w}_{j}$ and $f(x)$. However, $f()$ is not used to encode the context locations $\mathcal{C}^{t}$ before they are passed to $m_{\bm{\psi}}()$. Using these two different encoders performs increasingly poorly as the amount of context information increases.
@@
Figure A14: Impact of Location Encoder Training. Here we evaluate the performance of SINR and FS-SINR models when the size of the training dataset for the SINR backbone is varied. Results for FS-SINR models are shown with standard deviations from three runs (left), and without (right) for clarity. Evaluation on FS-SINR is performed with ‘Range Text’, while SINR can only make use of location data. ‘1000’, ‘100’, and ‘10’ represent the maximum number of examples per class the SINR backbone was trained on. ‘SINR (rand_init)’ is initialized with random weights and is not trained. ‘(trained on eval species)’ indicates training on all training and evaluation species.
@@
Figure A15: Impact of location encoder. Here we evaluate the performance of FS-SINR style models with different location encoders. Results are shown with standard deviations from three runs (left), and without (right) for clarity. Evaluation is performed with ‘Range Text’ on the IUCN dataset. ‘1000’, ‘100’, ‘10’ represent the maximum number of examples per class the SINR backbone was trained on. ‘(Frozen)’ indicates that the location encoder parameters were not updated during FS-SINR training. ‘FS-SATCLIP’ replaces the SINR location encoder with a pretrained, frozen location encoder from Klemmer et al. (2025). ‘FS-SINR (Fourier Location Encoder)’ uses the simple Fourier feature encoding (Tancik et al., 2020) used in Mildenhall et al. (2021) to match the 256 dimensional outputs of the SINR location encoders. These outputs are used directly as inputs to the transformer encoder. After a species token is produced in this way, it is attached to a pre-trained and fine-tuned SINR backbone to produce a range.
@@
Figure A16: Impact of training data. Here we evaluate FS-SINR models trained with different amounts of data. Results are shown with standard deviations from three runs (left), and without (right) for clarity. Evaluation is performed with ‘Range Text’ on the IUCN dataset. The labels show the maximum number of examples per-species that FS-SINR is trained on. We see that training on an intermediate amount of training data leads to best performance.
D.5 Ablating Training Data
In Figure A16 we vary the number of examples per-species that are provided during training. The impact of this is fairly small, with models trained on an intermediate amount of data performing best. It is worth noting, that not all species in the training dataset have as many as 1000 observations. We find that a model trained on only 10 examples per-species performs significantly worse.
@@ D.5 Ablating Training Data
Figure A17: Ablating model architecture components. Here we evaluate the performance of FS-SINR style models as we ablate various design choices. Results are shown with standard deviations from three runs (left), and without (right) for clarity. Evaluation is performed with ‘Range Text’ on the IUCN dataset. We see small changes in performance when removing the register token and the species decoder. However removing the learned token type embeddings has a large impact.
D.6 Ablating FS-SINR Architecture
In Figure A17 we vary the underlying FS-SINR architecture. Removing different components has a small effect on model performance, with the removal of the species decoder actually improving results when range text is provided. However, as several ablations perform very similarly, it is difficult to tease out the how much of this effect is due to variance. It is clear however that removing the learnable token type embeddings causes the model to completely fail to learn during training. In Figure A18 we show further ablations based around removing the learned location encoder for inputs to the transformer and replacing it with the simple Fourier feature encoding also seen in Figure A15. When this is removed, other ablations seem to further harm performance, although results for these ablations vary significantly between runs.
@@ D.6 Ablating FS-SINR Architecture
Figure A18: Further ablating model components. Here we evaluate the performance of FS-SINR style models as we ablate more components. Results are shown with standard deviations from three runs (left), and without (right) for clarity. Evaluation is performed with ‘Range Text’ on the IUCN dataset. ‘FF’ indicates that the model does not use a SINR backbone to encode location inputs to the transformer encoder. Instead, a simple Fourier feature encoding (Tancik et al., 2020) used in Mildenhall et al. (2021) is used to increase the dimensionality of location data to match the token dimension of the transformer encoder. These are used directly as inputs to the transformer encoder. After a species token is produced in this way, it is attached to a standard SINR backbone to produce a range. Removing the SINR backbone for encoding inputs to the transformer has a large impact on performance, especially when more context locations are supplied, and makes the model more sensitive to the impact of other ablations.
Appendix E Additional Qualitative Results
E.1 Visualizing Non-Species Concepts
@@
By jointly training on text and locations, FS-SINR is able to spatially ground abstract non-species concepts in a zero-shot manner, as is done with LE-SINR in Hamilton et al. (2024). In Figure A19 we provide another example similar to Figure 4 in the main paper. Here, we again fix the context location and show the impact of changing the text. We can see that different text prompts can result in quite different predicted ranges. In Figure A20 we see examples where different text concepts, which are very different from the species-based text provided during training, are grounded in sensible locations on the map. In Figure A21 we compare predictions made with increasing numbers of context locations in desert regions, with or without the accompanying text prompt “Desert”. As we increase the number of context locations, the two different models converge to more similar range predictions.
E.2 Visualizing Estimated Species Ranges
Here, we provide additional examples of the ranges produced by FS-SINR using context locations, text, and images. In Figures A22 and A23 we visualize FS-SINR range estimates for two different species when habitat or range text is provided. We observe that the combination of text and context locations seem to result in better estimates of the range. In Figure A24 we show additional zero-shot image-only examples, where FS-SINR is provided a single image from a held-out test species at inference time. Again, we observe some plausible range predictions even with such limited input data.
In Figure A26 we show range estimates for the Brown-banded Watersnake, using ‘range’ text for FS-SINR and LE-SINR approaches. In Figure A27 we show range estimates for the Brown-headed Honeyeater, using ‘habitat’ text for FS-SINR and LE-SINR approaches. Finally in Figure A28 we show range estimates for the Crevice Swift, without providing text. Overall, SINR produces more diffuse ranges and requires more locations to narrow down the range. LE-SINR and FS-SINR appear to have very different zero-shot behaviors, with LE-SINR frequently seeming to predict presence in almost no locations at all, while FS-SINR tends to produce a zero-shot range that is too large.
@@ E.2 Visualizing Estimated Species Ranges
In Figure A25 we visualize FS-SINR range predictions for the Yellow-footed Green Pigeon for models that have had different random initializations (i.e., different random seeds). We observe that there is a relatively large amount of variance in the outputs produced given the same input data. The same set of input context locations could represent many different possible output ranges, and thus being able to represent this variety is advantageous. We also utilize these different predictions from different seeds for ensembling and uncertainty quantification in Appendix A.
[…]
@@
European Robin - Range Text European Robin - Habitat Text
@@
Refer to caption Refer to caption
Range Text: The European robin is found across Europe, east to Western Siberia and south to North Africa; it is sedentary in most of its range except the far north. It also occurs in the Atlantic islands as far west as the Central Group of the Azores and Madeira. It is a vagrant in Iceland and has been introduced to other regions, including North America and Australia, but these introductions were unsuccessful. Habitat Text: The European robin inhabits a variety of habitats, including gardens, parks, woodlands, and forests. It prefers areas with dense vegetation and is often found near human settlements. It is also found in mountainous regions and can be seen in urban areas, such as cities and towns.
Figure A22: Using text descriptions. Here we illustrate the zero-shot (top row) and one-shot (bottom row) FS-SINR range estimations based on text descriptions for the European Robin, using ‘Range’ (left), and ‘Habitat’ (right) text, shown below the range estimates. Expert-derived range maps are shown inset in the top row.
American Pika - Range Text American Pika - Habitat Text
@@
Refer to caption Refer to caption
Range Text: The American pika (Ochotona princeps) is found in the mountains of western North America, usually in boulder fields at or above the tree line, from central British Columbia and Alberta in Canada to the US states of Oregon, Washington, Idaho, Montana, Wyoming, Colorado, Utah, Nevada, California, and New Mexico. Habitat Text: Pikas inhabit talus fields that are fringed by suitable vegetation in alpine areas. They also live in piles of broken rock. Sometimes, they live in man-made substrate such as mine tailings and piles of scrap lumber. Pikas usually have their den and nest sites below rock, around 20-100 cm (8-39 in) in diameter, but often sit on larger and more prominent rocks.
Figure A23: Using text descriptions. Here we illustrate the zero-shot (top row) and one-shot (bottom row) FS-SINR range estimations based on text descriptions for the American Pika, using ‘Range’ (left), and ‘Habitat’ (right) text, shown below the range estimates. Expert-derived range maps are shown inset.
Figure A24: Image zero-shot range estimation. Here we see zero-shot range estimates for six species in the IUCN evaluation dataset, with expert-derived range and image inset. The blue duck image taken from iNaturalist (2025) only shows evidence of the species from footprints in wet sand. We see that this image generates predictions in coastal areas in various locations around the globe. The coastal background for the Black Turnstone could have helped the model to generate a relatively accurate prediction on the northwest coast of North America.
@@ 20 Context
Figure A26: Comparing estimated ranges across models with context lcoations and range text. Here we see zero-shot and few-shot range estimates produced by FS-SINR, LE-SINR, and SINR for the Brown-banded Watersnake, with expert-derived range inset. We provide range text to FS-SINR and LE-SINR as well as context locations, but SINR is not capable of accepting text and so we show a blank map for the zero-shot range estimate. We see that LE-SINR underestimates the range using only text, while FS-SINR overestimates it. SINR requires more location data than the other approaches to localize the range to South America.
@@
Figure A27: Comparing estimated ranges across models with context locations and habitat text. Here we see zero-shot and few-shot range estimates produced by FS-SINR, LE-SINR, and SINR for the Brown-headed Honeyeater, with expert-derived range inset. We provide habitat text to FS-SINR and LE-SINR as well as context locations, but SINR is not capable of accepting text and so we show a blank map for the zero-shot range estimate. We again see LE-SINR underestimate the range using only text, while FS-SINR has very good zero-shot performance for this species. We see that SINR again requires more location data to narrow down the range and even after 20 locations the range is still significantly larger than the other models, extending into South Africa.